Assume that a sample is used to estimate a population proportion p. Find the margin of error E that corresponds to the given statistics and confidence level. Round the margin of error to four decimal places.

95% confidence; n = 2015, x = 1631
Answer #1

\hat{p}=\frac{x}{n}=\frac{1631}{2015}=0.8094

Confidence level is 95%

Margin of error is

E=1.96*\sqrt{\frac{p(1-p)}{n}}

E=1.96*\sqrt{\frac{0.8094(1-0.8094)}{2015}}=0.0172